# EHRAgent: Code Empowers Large Language Models for Complex Tabular Reasoning on Electronic Health Records

### v1

@@ Abstract
Large language models (LLMs) have demonstrated exceptional capabilities in planning and tool utilization as autonomous agents, but few have been developed for medical problem-solving. We propose EHRAgent  [^1] Our implementation of EHRAgent is available at https://github.com/wshi83/EhrAgent., an LLM agent empowered with a code interface, to autonomously generate and execute code for complex clinical tasks within electronic health records (EHRs). First, we formulate an EHR question-answering task into a tool-use planning process, efficiently decomposing a complicated task into a sequence of manageable actions. By integrating interactive coding and execution feedback, EHRAgent learns from error messages and improves the originally generated code through iterations. Furthermore, we enhance the LLM agent by incorporating long-term memory, which allows EHRAgent to effectively select and build upon the most relevant successful cases from past experiences. Experiments on two real-world EHR datasets show that EHRAgent outperforms the strongest LLM agent baseline by 36.48% and 12.41%, respectively. EHRAgent leverages the emerging few-shot learning capabilities of LLMs, enabling autonomous code generation and execution to tackle complex clinical tasks with minimal demonstrations.
EHRAgent: Code Empowers Large Language Models for Complex
Tabular Reasoning on Electronic Health Records
[…]
@@ 1 Introduction
Figure 1: Simple and efficient interactions between clinicians and EHR systems with the assistance of LLM agents. Clinicians specify tasks in natural language, and the LLM agent autonomously generates and executes code to interact with EHRs (right) for answers. It eliminates the need for specialized expertise or extra effort from data engineers, which is typically required when dealing with EHRs in clinical settings (left).
@@
An electronic health record (EHR) is a digital version of a patient’s medical history maintained by healthcare providers over time (Gunter and Terry, 2005). In clinical research and practice, clinicians actively interact with EHR systems to access and retrieve patient data, ranging from detailed individual-level records to comprehensive population-level insights (Cowie et al., 2017; Xu et al., 2022). Since most EHRs use pre-defined rule-based conversation systems (e.g., Epic), clinicians may require additional training or seek assistance from data engineers when obtaining information beyond the existing rules (Mandel et al., 2016; Bender and Sartipi, 2013). Alternatively, an autonomous agent could facilitate clinicians to communicate with EHRs in natural languages, translating clinical questions into machine-interpretable queries (Lee et al., 2022), planning a sequence of actions, and ultimately delivering the final responses, which holds great potential to simplify workflows and reduce workloads for clinicians (Figure 1).
@@
Large language models (LLMs) (OpenAI, 2023; Anil et al., 2023) bring us one step closer to autonomous agents, with extensive knowledge and substantial instruction-following abilities from diverse corpora during pretraining. LLM-based autonomous agents have demonstrated remarkable capabilities in problem-solving, such as reasoning (Wei et al., 2022; Wang et al., 2023d; Zhou et al., 2023), planning (Yao et al., 2023b; Liu et al., 2023; Sun et al., 2023; Hao et al., 2023), and memorizing (Wang et al., 2023b). One particularly notable capability of LLM agents is tool-usage (Schick et al., 2023; Qin et al., 2023a; Zhuang et al., 2023b), where they can utilize external tools (e.g., calculators, APIs, etc.), interact with environments, and generate action plans with intermediate reasoning steps that can be executed sequentially towards a valid solution (Wu et al., 2023b; Zhang et al., 2023).
@@
Despite their success in general-domain tasks, LLMs have encountered unique but significant challenges when it comes to real-world clinical research and practice (Jiang et al., 2023; Yang et al., 2022; Xu et al., 2023; Moor et al., 2023; Shi et al., 2023), especially for EHRs that have complex structures and require additional information and expertise beyond their pre-trained data. First, given the constraints in both the volume and specificity of training data within the medical field, LLMs still struggle with medical reasoning due to insufficient knowledge and understanding of EHRs (Thapa and Adhikari, 2023). Second, EHRs are typically relational databases containing vast amounts of tables (e.g., 26 tables in MIMIC-III (Johnson et al., 2016)) with heterogeneous patient data, including both administrative and clinical information. Moreover, unlike standardized questions (e.g., multi-choice) found in medical licensing exams (Jin et al., 2021), real-world clinical tasks are highly diverse and complex (Lee et al., 2022). These questions often arise from the unique circumstances of individual patients or specific groups, necessitating multi-step or complicated operations.
@@
To address these limitations, we propose EHRAgent, an autonomous LLM agent with external tools and code interface for improved medical reasoning across EHRs. We transform the EHR question-answering problem into a tool-use planning process: generating, executing, debugging, and optimizing a sequence of code-based actions. To overcome the lack of domain knowledge, we integrate additional information (e.g., detailed descriptions of each table in EHRs) and clinical knowledge by instructing the LLM agent to retrieve the most relevant knowledge. We then establish an interactive coding mechanism, which involves a multi-turn dialogue between the code planner and executor, iteratively refining the generated code-based plan. Specifically, EHRAgent optimizes the execution plan by incorporating environment feedback and delving further into error information to enhance debugging proficiency. In addition, we take advantage of long-term memory to continuously maintain a set of successful cases and dynamically select the most relevant few-shot examples, in order to effectively learn from and improve upon past experiences.
@@
We conducted extensive experiments on two widely used real-world EHR datasets, MIMIC-III (Johnson et al., 2016) and eICU (Pollard et al., 2018), to validate the empirical effectiveness of EHRAgent, with a particular focus on challenging tasks that align with real-world application scenarios. In contrast to traditional supervised learning methods that require extensive training samples with fine-grained annotations (e.g., text-to-SQL (Lee et al., 2022)), EHRAgent demonstrates its efficiency by necessitating only four demonstrations. Our findings suggest that EHRAgent enables complex tabular reasoning on EHRs by autonomous code generation and execution with environmental feedback. To the best of our knowledge, EHRAgent represents one of the first LLM agents for complex reasoning on EHRs with external tools and code interface.
@@
$\bullet$ We propose EHRAgent, an LLM agent augmented with tools and medical knowledge, to solve complex tabular reasoning derived from EHRs;
$\bullet$ Planning with a code interface, EHRAgent enables the LLM agent to formulate a clinical problem-solving process as an executable code plan of action sequences, along with a code executor;
$\bullet$ We introduce interactive coding between the LLM agent and code executor, iteratively refining plan generation and optimizing code execution by examining environment feedback in depth.
Figure 2: Overview of our proposed LLM agent, EHRAgent, for complex tabular reasoning tasks on EHRs. Given an input clinical question based on EHRs, EHRAgent initially incorporates relevant medical knowledge. Subsequently, EHRAgent decomposes the task and generates a plan (i.e., code) based on EHR descriptions, tool function definitions, few-shot examples, and integrated medical knowledge. Upon execution, EHRAgent iteratively debugs the code following the environmental feedback and ultimately generates the final solution.
@@ 2 Preliminaries
LLM Agent Setup. We further formulate the planning process for LLMs as autonomous agents in EHR question answering. For initialization, the LLM agent is equipped with a set of pre-built tools =$\mathcal{M}=\{M_{0},M_{1},\cdots\}$ to interact with and address queries derived from EHRs $\mathcal{R}$. Given an input query $q\in\mathcal{Q}$ from the task space $\mathcal{Q}$, the objective of the LLM agent is to design a $T$-step execution plan $P=(a_{1},a_{2},\cdots,a_{T})$, with each action $a_{t}$ selected from the tool set $a_{t}\in\mathcal{M}$. Specifically, we generate the action sequences (i.e., plan) by prompting the LLM agent following a policy $p_{q}\sim\pi(a_{0},a_{1},\cdots,a_{T_{q}}|q;\mathcal{R},\mathcal{M}):\mathcal{Q}\times\mathcal{R}\times\mathcal{M}\to\Delta(\mathcal{M})^{T_{q}}$, where $\Delta(\cdot)$ is a probability simplex function. The final output is obtained by executing the entire plan $y\sim\rho(y|q,a_{0},a_{1},\cdots,a_{T_{q}})$, where $\rho$ is a plan executor interacting with EHRs.
Planning with Code Interface. To mitigate ambiguities and misinterpretations in plan generation, an increasing number of LLM agents Gao et al. (2023); Liang et al. (2023); Sun et al. (2023); Chen et al. (2023a); Zhuang et al. (2023a) employ code as planner interface instead of natural language for prompting. The code interface enables LLM agents to formulate an executable code plan as action sequences, intuitively transforming natural language question-answering into iterative coding Yang et al. (2023). Consequently, the planning policy $\pi(\cdot)$ turns into a code generation process, with a code execution as the executor $\rho(\cdot)$. We then track the outcome of each interaction back to the LLM agent, which can be either a successful execution result or an error message, to iteratively refine the generated code-based plan. This interactive process, a multi-turn dialogue between the planner and executor, takes advantage of the advanced reasoning capabilities of LLMs to optimize plan refinement and execution.
@@ 3 EHRAgent: LLMs as Medical Agents
In this section, we present EHRAgent (Figure 2), an LLM agent that enables multi-turn interactive coding to address multi-hop reasoning tasks on EHRs. EHRAgent comprises four key components: (1) Medical Knowledge Integration: EHRAgent summarizes the most important relevant information to facilitate a comprehensive understanding of EHRs. (2) Interactive Coding with Execution Feedback: EHRAgent harnesses LLMs as assistant agents in a multi-turn conversation with a code executor. (3) Debugging via Error Information Tracing: Rather than simply sending back information from the code executor, EHRAgent thoroughly analyzes error messages to identify the root causes through iterations until a final solution. (4) Plan Refinement with Long-Term Memory: Using long-term memory, EHRAgent selects the most relevant successful cases as demonstrations from past experiences for effective plan refinement. We summarize the workflow of EHRAgent in Algorithm 1.
Input: $q$: input question; $\mathcal{R}$: reference EHRs; $\mathcal{C}_{i}$: column description of EHR $R_{i}$; $\mathcal{D}$: descriptions of EHRs $\mathcal{R}$; $T$: the maximum number of steps; $\mathcal{T}$: definitions of tool function.
@@
// Medical Knowledge Integration
@@
Algorithm 1 The procedure of EHRAgent.
3.1 Medical Knowledge Integration
We first incorporate medical knowledge into EHRAgent for a comprehensive understanding of EHRs within a limited context length. Given an EHR-based clinical question $q$ and the reference EHRs =$\mathcal{R}=\{R_{0},R_{1},\cdots\}$, the objective of knowledge integration is to generate descriptions of knowledge most relevant to $q$, thereby facilitating the identification and location of potential useful references within $\mathcal{R}$. For example, given a query related to ‘Aspirin’, we expect LLMs to locate the drug ‘Aspirin’ at the PRESCRIPTION table, under the prescription_name column in the EHR.
To achieve this, we initially maintain a thorough introduction $\mathcal{I}$ of all the reference EHRs, including overall data descriptions $\mathcal{D}$ and the detailed columnar descriptions $\mathcal{C}_{i}$ for each individual EHR $R_{i}$, expressed as $\mathcal{I}=[\mathcal{D};\mathcal{C}_{0};\mathcal{C}_{1};\cdots]$. To further extract additional background knowledge essential for addressing the complex query $q$, we then distill key information from the detailed introduction $\mathcal{I}$. Specifically, we directly prompt LLMs to generate the relevant knowledge $B(q)$ based on few-shot demonstrations, denoted as $B(q)=\operatorname{LLM}([\mathcal{I};q])$.
@@ 3.2 Interactive Coding with Execution
We then introduce interactive coding between the LLM agent (i.e., code generator) and code executor to facilitate iterative plan refinement. EHRAgent integrates LLMs as an assistant agent with a code executor within a multi-turn conversation. The code executor retrieves and executes the generated code and then provides the execution results back to the LLM. Within the conversation, EHRAgent navigates the subsequent phase of the dialogue, where the LLM agent is expected to either (1) continue to iteratively adjust its original code in response to any errors encountered or (2) finally deliver a conclusive answer based on the successful execution outcomes.
LLM Agent. To generate accurate code snippets $C(q)$ as solution plans for the query $q$, we prompt the LLM agent with a combination of the EHR introduction $\mathcal{I}$, tool function definitions $\mathcal{T}$, a set of $K$-shot examples $E_{1},\cdots,E_{K}$, the input query $q$, and the integrated medical knowledge relevant to the query $B(q)$:
$C(q)=\operatorname{LLM}([\mathcal{I};\mathcal{T};E_{1},\cdots,E_{K};q;B(q)]).$ (1)
Leveraging the AutoGen infrastructure (Wu et al., 2023b) of automated multi-agent conversation, we develop the LLM agent to (1) generate code within a designated coding block as required, (2) modify the code according to the outcomes of its execution, and (3) insert a specific code “TERMINATE” at the end of its response to indicate the conclusion of the conversation.
Code Executor. The code executor automatically extracts the code from the LLM agent’s output and executes it within the local environment: $O(q)=\operatorname{EXECUTE}(C(q))$. After execution, it sends back the execution results to the LLM agent for potential plan refinement and further processing.
3.3 Error Tracing via Rubber Duck Debugging
Our empirical observations indicate that LLM agents tend to make slight modifications to the code snippets based on the error message without further debugging. In contrast, human programmers often delve deeper, identifying bugs or underlying causes by analyzing the code implementation against the error descriptions (Chen et al., 2023b). Inspired by this, we apply a ‘rubber duck debugging’ pipeline for plan refinement with the LLM agent. Specifically, we provide detailed trace feedback, including error type, message, and location, all parsed from the error information by the code executor. Subsequently, this error context is presented to a ‘rubber duck’ LLM, prompting it to trace the most probable causes of the error. The generated explanations are then fed back into the conversation flow, aiding in the debugging process. For the $t$-th interaction between the LLM agent and the code executor, the process is as follows:

### v3

# EHRAgent: Code Empowers Large Language Models for Few-shot Complex Tabular Reasoning on Electronic Health Records

@@ Abstract
Clinicians often rely on data engineers to retrieve complex patient information from electronic health record (EHR) systems, a process that is both inefficient and time-consuming. We propose EHRAgent  [^1] Our implementation of EHRAgent is available at https://github.com/wshi83/EhrAgent., a large language model (LLM) agent empowered with accumulative domain knowledge and robust coding capability. EHRAgent enables autonomous code generation and execution to facilitate clinicians in directly interacting with EHRs using natural language. Specifically, we formulate a multi-tabular reasoning task based on EHRs as a tool-use planning process, efficiently decomposing a complex task into a sequence of manageable actions with external toolsets. We first inject relevant medical information to enable EHRAgent to effectively reason about the given query, identifying and extracting the required records from the appropriate tables. By integrating interactive coding and execution feedback, EHRAgent then effectively learns from error messages and iteratively improves its originally generated code. Experiments on three real-world EHR datasets show that EHRAgent outperforms the strongest baseline by up to 29.6% in success rate, verifying its strong capacity to tackle complex clinical tasks with minimal demonstrations.
EHRAgent: Code Empowers Large Language Models for Few-shot
Complex Tabular Reasoning on Electronic Health Records
Wenqi Shi1thanks:   Equal contribution.  Ran Xu211footnotemark: 1  Yuchen Zhuang1  Yue Yu1  Jieyu Zhang3  Hang Wu1  Yuanda Zhu1  Joyce Ho2  Carl Yang2  May D. Wang1 1 Georgia Institute of Technology  2 Emory University  3 University of Washington {wqshi,yczhuang,yueyu,hangwu,yzhu94,maywang}@gatech.edu, {ran.xu,joyce.c.ho,j.carlyang}@emory.edu,  jieyuz2@cs.washington.edu
@@ 1 Introduction
Figure 1: Simple and efficient interactions between clinicians and EHR systems with the assistance of LLM agents. Clinicians specify tasks in natural language, and the LLM agent autonomously generates and executes code to interact with EHRs (right) for answers. It eliminates the need for specialized expertise or extra effort from data engineers, which is typically required when dealing with EHRs in existing clinical settings (left).
@@
An electronic health record (EHR) is a digital version of a patient’s medical history maintained by healthcare providers over time (Gunter and Terry, 2005). In clinical research and practice, clinicians actively interact with EHR systems to access and retrieve patient data, ranging from detailed individual-level records to comprehensive population-level insights (Cowie et al., 2017). The reliance on pre-defined rule-based conversion systems in most EHRs often necessitates additional training or assistance from data engineers for clinicians to obtain information beyond these rules (Mandel et al., 2016; Bender and Sartipi, 2013), leading to inefficiencies and delays that may impact the quality and timeliness of patient care.
Alternatively, an autonomous agent could facilitate clinicians to communicate with EHRs in natural languages, translating clinical questions into machine-interpretable queries, planning a sequence of actions, and ultimately delivering the final responses. Compared to existing EHR management that relies heavily on human effort, the adoption of autonomous agents holds great potential to efficiently simplify workflows and reduce workloads for clinicians (Figure 1). Although several supervised learning approaches (Lee et al., 2022; Wang et al., 2020) have been explored to automate the translation of clinical questions into corresponding machine queries, such systems require extensive training samples with fine-grained annotations, which are both expensive and challenging to obtain.
@@
Large language models (LLMs) (OpenAI, 2023; Anil et al., 2023) bring us one step closer to autonomous agents with extensive knowledge and substantial instruction-following abilities from diverse corpora during pretraining. LLM-based autonomous agents have demonstrated remarkable capabilities in problem-solving, such as reasoning (Wei et al., 2022), planning (Yao et al., 2023b), and memorizing (Wang et al., 2023b). One particularly notable capability of LLM agents is tool-usage (Schick et al., 2023; Qin et al., 2023), where they can utilize external tools (e.g., calculators, APIs, etc.), interact with environments, and generate action plans with intermediate reasoning steps that can be executed sequentially towards a valid solution (Wu et al., 2023; Zhang et al., 2023).
(a) # Rows(k) per Table
(b) # Tables per Question
Figure 2: Compared to general domain tasks (blue) such as WikiSQL (Zhong et al., 2017) and SPIDER (Yu et al., 2018), multi-tabular reasoning tasks within EHRs (orange) typically involve a significantly larger number of records per table and necessitate querying multiple tables to answer each question, thereby requiring more advanced reasoning and problem-solving capabilities.
@@
Despite their success in general domains, LLMs have encountered unique and significant challenges in the medical domain (Jiang et al., 2023; Yang et al., 2022; Moor et al., 2023), especially when dealing with individual EHR queries that require advanced reasoning across a vast number of records within multiple tables (Li et al., 2024; Lee et al., 2022) (Figure 2). First, given the constraints in both the volume and specificity of training data within the medical field (Thapa and Adhikari, 2023), LLMs still struggle to identify and extract relevant information from the appropriate tables and records within EHRs, due to insufficient knowledge and understanding of their complex structure and content. Second, EHRs are typically large-scale relational databases containing vast amounts of tables with comprehensive administrative and clinical information (e.g., 26 tables of 46K patients in MIMIC-III). Moreover, real-world clinical tasks derived from individual patients or specific groups are highly diverse and complex, requiring multi-step or complicated operations.
@@
To address these limitations, we propose EHRAgent, an autonomous LLM agent with external tools and code interface for improved multi-tabular reasoning across EHRs. We translate the EHR question-answering problem into a tool-use planning process – generating, executing, debugging, and optimizing a sequence of code-based actions. Firstly, to overcome the lack of domain knowledge in LLMs, we instruct EHRAgent to integrate query-specific medical information for effectively reasoning from the given query and locating the query-related tables or records. Moreover, we incorporate long-term memory to continuously maintain a set of successful cases and dynamically select the most relevant few-shot examples, in order to effectively learn from and improve upon past experiences. Secondly, we establish an interactive coding mechanism, which involves a multi-turn dialogue between the code planner and executor, iteratively refining the generated code-based plan for complex multi-hop reasoning. Specifically, EHRAgent optimizes the execution plan by incorporating environment feedback and delving into error messages to enhance debugging proficiency.
@@
We conduct extensive experiments on three large-scale real-world EHR datasets to validate the empirical effectiveness of EHRAgent, with a particular focus on challenging tasks that reflect diverse information needs and align with real-world application scenarios. In contrast to traditional supervised settings Lee et al. (2022); Wang et al. (2020) that require over 10K training samples with manually crafted annotations, EHRAgent demonstrates its efficiency by necessitating only four demonstrations. Our findings suggest that EHRAgent improves multi-tabular reasoning on EHRs through autonomous code generation and execution, leveraging accumulative domain knowledge and interactive environmental feedback.
@@
$\bullet$ We propose EHRAgent, an LLM agent augmented with external tools and domain knowledge, to solve few-shot multi-tabular reasoning derived from EHRs with only four demonstrations;
$\bullet$ Planning with a code interface, EHRAgent formulates a complex clinical problem-solving process as an executable code plan of action sequences, along with a code executor;
$\bullet$ We introduce interactive coding between the LLM agent and code executor, iteratively refining plan generation and optimizing code execution by examining environmental feedback in depth;
$\bullet$ Experiments on three EHR datasets show that EHRAgent improves the strongest baseline on multi-hop reasoning by up to 29.6% in success rate.
Figure 3: Overview of our proposed LLM agent, EHRAgent, for complex few-shot tabular reasoning tasks on EHRs. Given an input clinical question based on EHRs, EHRAgent decomposes the task and generates a plan (i.e., code) based on (a) metadata (i.e., descriptions of tables and columns in EHRs), (b) tool function definitions, (c) few-shot examples, and (d) domain knowledge (i.e., integrated medical information). Upon execution, EHRAgent iteratively debugs the generated code following the execution errors and ultimately generates the final solution.
@@ 2 Preliminaries
LLM Agent Setup. We further formulate the planning process for LLMs as autonomous agents in EHR question answering. For initialization, the LLM agent is equipped with a set of pre-built tools =$\mathcal{M}=\{M_{0},M_{1},\cdots\}$ to interact with and address queries derived from EHRs $\mathcal{R}$. Given an input query $q\in\mathcal{Q}$ from the task space $\mathcal{Q}$, the objective of the LLM agent is to design a $T$-step execution plan $P=(a_{1},a_{2},\cdots,a_{T})$, with each action $a_{t}$ selected from the tool set $a_{t}\in\mathcal{M}$. Specifically, we generate the action sequences (i.e., plan) by prompting the LLM agent following a policy $p_{q}\sim\pi(a_{1},\cdots,a_{T_{q}}|q;\mathcal{R},\mathcal{M}):\mathcal{Q}\times\mathcal{R}\times\mathcal{M}\to\Delta(\mathcal{M})^{T_{q}}$, where $\Delta(\cdot)$ is a probability simplex function. The final output is obtained by executing the entire plan $y\sim\rho(y|q,a_{1},\cdots,a_{T_{q}})$, where $\rho$ is a plan executor interacting with EHRs.
Planning with Code Interface. To mitigate ambiguities and misinterpretations in plan generation, an increasing number of LLM agents (Gao et al., 2023; Liang et al., 2023; Sun et al., 2023; Chen et al., 2023; Zhuang et al., 2024) employ code prompts as planner interface instead of natural language prompts. The code interface enables LLM agents to formulate an executable code plan as action sequences, intuitively transforming natural language question-answering into iterative coding (Yang et al., 2023). Consequently, the planning policy $\pi(\cdot)$ turns into a code generation process, with a code execution as the executor $\rho(\cdot)$. We then track the outcome of each interaction back to the LLM agent, which can be either a successful execution result or an error message, to iteratively refine the generated code-based plan. This interactive process, a multi-turn dialogue between the planner and executor, takes advantage of the advanced reasoning capabilities of LLMs to optimize plan refinement and execution.
@@ 3 EHRAgent: LLMs as Medical Agents
In this section, we present EHRAgent (Figure 3), an LLM agent that enables multi-turn interactive coding to address multi-hop reasoning tasks on EHRs. EHRAgent comprises four key components: (1) Medical Information Integration: We incorporate query-specific medical information for effective reasoning based on the given query, enabling EHRAgent to identify and retrieve the necessary tables and records for answering the question. (2) Demonstration Optimization through Long-Term Memory: Using long-term memory, EHRAgent replaces original few-shot demonstrations with the most relevant successful cases retrieved from past experiences. (3) Interactive Coding with Execution Feedback: EHRAgent harnesses LLMs as autonomous agents in a multi-turn conversation with a code executor. (4) Rubber Duck Debugging via Error Tracing: Rather than simply sending back information from the code executor, EHRAgent thoroughly analyzes error messages to identify the underlying causes of errors through iterations until a final solution. We summarize the workflow of EHRAgent in Algorithm 1.
Input: $q$: input question; $\mathcal{R}$: reference EHRs; $\mathcal{C}_{i}$: column description of EHR $R_{i}$; $\mathcal{D}$: descriptions of EHRs $\mathcal{R}$; $T$: the maximum number of steps; $\mathcal{T}$: definitions of tool function; $\mathcal{L}$: long-term memory.
@@
// Medical Information Integration
@@
Algorithm 1 Overview of EHRAgent.
3.1 Medical Information Integration
Clinicians frequently pose complex inquiries that necessitate advanced reasoning across multiple tables and access to a vast number of records within a single query. To accurately identify the required tables, we first incorporate query-specific medical information (i.e., domain knowledge) into EHRAgent to develop a comprehensive understanding of the query within a limited context length. Given an EHR-based clinical question $q$ and the reference EHRs =$\mathcal{R}=\{R_{0},R_{1},\cdots\}$, the objective of information integration is to generate the domain knowledge most relevant to $q$, thereby facilitating the identification and location of potential useful references within $\mathcal{R}$. For example, given a query related to ‘Aspirin’, we expect LLMs to locate the drug ‘Aspirin’ at the PRESCRIPTION table, under the prescription_name column in the EHR.
To achieve this, we initially maintain a thorough metadata $\mathcal{I}$ of all the reference EHRs, including overall data descriptions $\mathcal{D}$ and the detailed column descriptions $\mathcal{C}_{i}$ for each individual EHR $R_{i}$, expressed as $\mathcal{I}=[\mathcal{D};\mathcal{C}_{0};\mathcal{C}_{1};\cdots]$. To further extract additional background knowledge essential for addressing the complex query $q$, we then distill key information from the detailed introduction $\mathcal{I}$. Specifically, we directly prompt LLMs to generate the relevant information $B(q)$ based on demonstrations, denoted as $B(q)=\operatorname{LLM}([\mathcal{I};q])$.
3.2 Demonstration Optimization through Long-Term Memory
where $\arg\operatorname{TopK}{\max}(\cdot)$ identifies the indices of the top $K$ elements with the highest values from $\mathcal{L}$, and $\operatorname{sim}(\cdot,\cdot)$ calculates the similarity between two questions, employing negative Levenshtein distance as the similarity metric. Following this retrieval process, the newly acquired $K$-shot examples $\mathcal{E}(q)$ replace the originally predefined examples =$\mathcal{E}=\{E_{1},\cdots,E_{K}\}$. This updated set of examples serves to reformulate the prompt, guiding EHRAgent in optimal demonstration selection by leveraging accumulative domain knowledge.
@@ 3.3 Interactive Coding with Execution
We then introduce interactive coding between the LLM agent (i.e., code generator) and code executor to facilitate iterative plan refinement. EHRAgent integrates LLMs with a code executor in a multi-turn conversation. The code executor runs the generated code and returns the results to the LLM. Within the conversation, EHRAgent navigates the subsequent phase of the dialogue, where the LLM agent is expected to either (1) continue to iteratively refine its original code in response to any errors encountered or (2) finally deliver a conclusive answer based on the successful execution outcomes.
LLM Agent. To generate accurate code snippets $C(q)$ as solution plans for the query $q$, we prompt the LLM agent with a combination of the EHR introduction $\mathcal{I}$, tool function definitions $\mathcal{T}$, a set of $K$-shot examples $\mathcal{E}(q)$ updated by long-term memory, the input query $q$, and the integrated medical information relevant to the query $B(q)$:
We develop the LLM agent to (1) generate code within a designated coding block as required, (2) modify the code according to the outcomes of its execution, and (3) insert a specific code “TERMINATE” at the end of its response to indicate the conclusion of the conversation.
Code Executor. The code executor automatically extracts the code from the LLM agent’s output and executes it within the local environment: $O(q)=\operatorname{EXECUTE}(C(q))$. After execution, it sends back the execution results to the LLM agent for potential plan refinement and further processing. Given the alignment of empirical observations and Python’s inherent modularity with tool functions [^2] We include additional analysis in Appendix D to further justify the selection of primary programming language., we select Python 3.9 as the primary coding language for interactions between the LLM agent and the code executor.
3.4 Rubber Duck Debugging via Error Tracing
Our empirical observations indicate that LLM agents tend to make slight modifications to the code snippets based on the error message without further debugging. In contrast, human programmers often delve deeper, identifying bugs or underlying causes by analyzing the code implementation against the error descriptions (Chen et al., 2024). Inspired by this, we integrate a ‘rubber duck debugging’ pipeline with error tracing to refine plans with the LLM agent. Specifically, we provide detailed trace feedback, including error type, message, and location, all parsed from the error information by the code executor. Subsequently, this error context is presented to a ‘rubber duck’ LLM, prompting it to generate the most probable causes of the error. The generated explanations are then fed back into the conversation flow, aiding in the debugging process. For the $t$-th interaction between the LLM agent and the code executor, the process is as follows: